Here are some new plots of the 100MHz transmission PD. I removed the 200MHz notch that I had put in (since we don't need it), and saw that the resonance at high frequency looked pretty bad. I wanted to check if this was because of the low gain we were using (datasheet says that it is stable for gain>=10), so I retook the TF and then changed the gain up to 3 and then 10 by changing R5 to 100 and then 453.
The gain3 is an improvement, but there's still some nastiness at high freqency. The gain 10 plot looks pretty clean though. Compared to the liso model (which I have set to gain 10, and have changed the diode capacitance down to 100pF) the gain seems to be a lot higher than we should be getting. I've gone over the calibration a few times and don't see any errors. I've also checked the values in liso and they seem correct too. This is not great since we really don't want to have such a high gain in the first place.
We should probably see if there is a low gain version of the LMH6624 that we can use in its place.
I should also mention that I did take into account the difference in the laser power on the two PDs. The 1611 had 1.16mW on it, and our PD had 0.47mW on it.